Circular No. 8956

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V2670 OPHIUCHI R. W. Russell, R. J. Rudy, D. K. Lynch, and A. M. Gilbert, The

Aerospace Corporation; and C. E. Woodward, University of Minnesota, report 0.8- to 5.5-μm spectroscopy of V2670 Oph (cf. IAUC 8947, 8948, 8949) on June 14 UT using SpeX at the Infrared Telescope Facility. V2670 Oph (Nova Oph 2008 No. 1) displays a generally low-excitation spectrum with lines of neutral carbon, nitrogen, oxygen, and sodium, and singly ionized iron and calcium. V2670 Oph is probably an “Fe II”-type nova; it has distinct, tiered line profiles comprised of both narrow (FWHM = 1100 km/s) and broad (FWHM = 3000 km/s) components. The O I lines that are fluorescently excited by Lyβ are the strongest lines in the infrared spectrum and indicate a reddening of E(B −V) ≈ 1.3. There is no evidence at present for dust formation.

COMETS C/2008 E7–E9, C/2008 G3, AND C/2008 G4 (SOHO) Further to IAUC 8955, additional apparent comets have been found

on SOHO website images; all are Kreutz sungrazers except for C/2008 E7 (non-group). C/2008 E8 was small, stellar in appearance, and of mag ∼ 7. C/2008 E7 was visible in C3 images for a day and a half (peaking at mag ∼ 6 around Mar. 5.33 UT) before cutting across the C2 field-of-view and fading out; it was basically stellar in appearance, though it may have been slightly diffuse and marginally elongated (when it was very faint) in C2 images. C/2008 E9 was very faint (mag ∼ 7–7.5) and slightly diffuse in C3 images, and very diffuse and elongated in C2 images. C/2008 G3 was very diffuse and very faint (mag ∼ 8). C/2008 G4 was small, condensed, and teardrop-shaped (mag ∼ 7).

Comet 2008 UT α_{2000} δ_{2000} Inst. F MPEC
C/2008 E7 Mar. 4.638 22^{h}56.^{m}0 − 4^{o}27′ C3/2 HS 2008-L26
C/2008 E8 7.071 23 30.0 − 6 03 C3* HS 2008-L27
C/2008 E9  12.029 23 45.8 − 3 58 C3/2* BZ 2008-L27
C/2008 G3 Apr. 1.979 0 55.3 + 4 20 C2  HS 2008-L27
C/2008 G4  9.493 1 22.6 + 7 00 C2  MK 2008-L27

COMET 200P/LARSEN Comet P/2008 L1 = 1997 V1 (cf. IAUC 8952) has been assigned the

number 200P (cf. MPC 63085).

2008 June 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green